OpenStudy (anonymous):

What percent of 16 is 11?

OpenStudy (anonymous):

11/16... divide...

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

0.6875 :-)

OpenStudy (anonymous):

I did what Badref said... not it...

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

.6875

OpenStudy (anonymous):

Do you have multiple choices?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=11%2F16 Proof.

OpenStudy (anonymous):

No, briz.

OpenStudy (anonymous):

16x% = 11 x% = 11/16 = 68.75 %

OpenStudy (anonymous):

You got the right answer if it's 0.6875

OpenStudy (anonymous):

Maybe my program has a glitch then. It says that is not correct.

OpenStudy (anonymous):

Maybe :-)

OpenStudy (anonymous):

Here's a better proof, Abby. http://www.wolframalpha.com/input/?i=what+percent+of+16+is+11

OpenStudy (anonymous):

Your program is incorrect. My college math major and the world's best commercial calculator say 11/16=0.6875

OpenStudy (anonymous):

You just have to move the decimal over two spaces.

OpenStudy (anonymous):

That's true, but 0.68(1) is 68% of a unit value, so the multiplication is not entirely necessary.